OpenStudy (anonymous):

Can someone check my Algebra 2 answers please? I fan and medal!! 1. What is the simplified form of (x+1/x^2+x-6)/(x^2+5x+4/x-2) Answer: 1/(x+4)(x-2) 2. What is the simplified form of (2/x^2+x)-(1/x) Answer: 3-x/x(x-1) 3. What is the simplified form of (3/2x-5)+(21/8x^2-14x-15) Answer: 6(2x+5)/(2x-5)(4x+3) 4. What is the simplified form of (3/x^2)/(1/x^3) Answer: 1/3x 5. What is the simplified form of (15xy^2/x^2+5x+6)/(5x^2y/2x^2+7x+3) Answer: x(x-2)/3y(2x-1) 6. What are the discountinuities of the function f(x)=(x^2+5x+6)/(2x+16) Answer: x does not equal -8

OpenStudy (anonymous):

7. What are the vertical asymptotes of the function f(x)=(4x+8)/(x^2+3x-4) Answer: x=-1 and x=-4

OpenStudy (aum):

I am getting a different answer for 1.

OpenStudy (anonymous):

What did you get for number one?

OpenStudy (aum):

\[\frac{1}{(x+3)(x+4)}\]. Getting a different answer for #2 also. I don't want to end up giving all the answers.

OpenStudy (anonymous):

Okay I understand. Could you explain how you did number one?

OpenStudy (aum):

\[ \frac{ x+1}{x^2+x-6} \div \frac{x^2+5x+4}{x-2} = \\ \frac{(x+1)}{(x+3)(x-2)} \div \frac{(x+4)(x+1)}{(x-2)} = \\ \frac{(x+1)}{(x+3)(x-2)} \times \frac{(x-2)}{(x+4)(x+1)} = \\ \frac{\cancel{(x+1)}}{(x+3)\cancel{(x-2)}} \times \frac{\cancel{(x-2)}}{(x+4)\cancel{(x+1)}} = \frac{1}{(x+3)(x+4)}\\ \]

OpenStudy (anonymous):

Okay I see what I did wrong. Thank you!

OpenStudy (anonymous):

I redid number 2. This time I got (1-x)/x(x+1) is that correct?

OpenStudy (aum):

Yes.